# A Controllable and Realistic Framework for Evaluating Microservice Scheduling in Cloud-Edge Continuum

### v2

@@
Abstract.
The transition from traditional architectures to containerized microservices within the cloud-edge computing continuum introduces significant challenges, particularly in the efficient scheduling of microservices under dynamic conditions. Complex and fluctuating call-graph dependencies, varying cross-node communication latencies, and unpredictable bandwidth conditions substantially impact the performance and reliability of deployed microservices. Consequently, accurately evaluating scheduling policies in such dynamic environments remains essential yet challenging due to the lack of realistic and controllable evaluation frameworks.
@@ Abstract.
In this paper, we propose iDynamics, a novel evaluation framework designed explicitly to address these challenges. iDynamics provides comprehensive and controllable evaluation capabilities by emulating realistic dynamics, including configurable call-graph topologies, cross-node communication delays, and bandwidth variability. The framework is composed of modular components, such as the Graph Dynamics Analyzer, Networking Dynamics Manager, and Scheduling Policy Extender, enabling fine-grained environmental control and facilitating systematic comparisons of different scheduling strategies. Extensive experiments on a real cloud-edge testbed demonstrate that iDynamics effectively captures diverse dynamic scenarios encountered in microservice deployments, offering a robust solution for evaluating and optimizing policy performance under realistic and controllable conditions.
@@
Microservices, Policy Evaluation, Controllable Dynamics, Cloud-Edge Continuum
copyright: noneconference: ACM Conference; Month YYYY; City, Countryccs: Computer systems organization Cloud computingccs: Networks Network dynamicsccs: Software and its engineering Scheduling
1. Introduction
@@
Microservices have emerged as an important architecture in the cloud-edge computing continuum, revolutionizing the design and implementation of distributed applications. Characterized by their modularity and decentralization, microservices enhance flexibility and scalability, effectively meeting dynamic application demands (Pallewatta et al., 2023; Cerny et al., 2018; Boucher et al., 2018). However, this architectural transition introduces significant challenges, especially for managing complex workflows and intricate service dependencies (Gu et al., 2021; Luo et al., 2024; Pallewatta et al., 2024; Xu et al., 2022a). Unpredictable user requests, varying execution paths across services, and fluctuating network conditions exacerbate the difficulty of ensuring SLA (Service Level Agreement) compliance for applications deployed across cluster nodes.
@@
Within the cloud-edge continuum, numerous microservices interact frequently and are often shared among multiple applications (Liu et al., 2023; Luo et al., 2022). Such interactions can lead to performance degradation due to resource contention and cascading delays along request execution paths, directly impacting critical SLA-related metrics such as latency and throughput.
Given the complexities arising from workload variability, dynamic application call-graphs, and network-level fluctuations, fine-grained scheduling strategies and robust policy evaluation frameworks are essential. A well-designed policy evaluation framework would enable cloud providers to more accurately and efficiently accommodate diverse microservice applications while adhering to different SLA requirements.
@@
Although various frameworks have been proposed for evaluating scheduling policies for cloud microservices, existing solutions still have limitations. For instance, AutoARMOR (Li et al., 2021) automates inter-service access control policy generation but overlooks dynamic networking conditions and is limited in supporting diverse SLA evaluations. OptTraffic (Zhu et al., 2023) optimizes containerized microservice traffic but neglects cross-node latency fluctuations and introduces complexity due to pairwise analysis of dependent container replicas. Similarly, NetMARKS (Wojciechowski et al., 2021) utilizes dynamic network metrics from Istio Service Mesh for pod scheduling, yet it lacks comprehensive bi-directional traffic analysis and can lead to imbalanced load distributions. Other existing methods (Sun et al., 2024; Song et al., 2023; Cao and Sharma, 2021; Marchese and Tomarchio, 2022b, a) also face limitations regarding handling call-graph dynamics and networking variations.
@@
To address these gaps, we propose iDynamics, a novel framework that systematically manages multiple layers of dynamics, enabling precise evaluation of scheduling policies under realistic conditions. With iDynamics, users can observe the impact of dynamics on application performance and identify SLA violations effectively. Figure 4 illustrates the high-level design of our proposed framework, highlighting its modular components. To our knowledge, iDynamics is the first unified evaluation framework specifically designed with controllable dynamics for microservice applications in cloud-edge continuum environments. Our primary contributions are:
@@
The remainder of this paper is structured as follows: Section 2 reviews related work. Section 3 presents the background and motivation. Section 4 introduces the iDynamics framework and its components. Section 5 describes the design of the Graph Dynamics Analyzer, and Section 6 details the Networking Dynamics Manager. Section 7 explains the Scheduling Policy Extender. Section 8 provides performance evaluations and detailed case studies. Finally, Section 9 concludes the paper and outlines future directions.
2. Related Work
@@
The management of microservices in different environments has been extensively explored. This section reviews existing literature with an emphasis on microservice management, call-graph analysis, network dynamics emulation, and network-aware scheduling.
2.1. Microservice Management
@@
FIRM (Qiu et al., 2020) leverages telemetry metrics and machine learning models to fine-tune microservice management by identifying SLO violations and resource contentions, subsequently mitigating violations through reprovisioning actions. Erms (Luo et al., 2022) formulates resource scaling models specifically addressing latency goals for shared microservices with extensive call graphs. GrandSLAm (Kannan et al., 2019) predicts microservice execution stages’ completion times, dynamically batching and reordering requests based on these predictions. Nevertheless, these approaches often neglect dynamic cross-node delays and the impacts of bi-directional communication between replicated microservices.
2.2. Call-graph Dynamics Analysis
@@
Sage (Gan et al., 2021) employs graphical Bayesian models for root cause analysis of cascading QoS violations in interactive microservices, prioritizing practicality and scalability. Tian et al. (Tian et al., 2019) develop a synthetic workload generator for DAG jobs by analyzing large-scale cluster traces. Luo et al. (Luo et al., 2021) characterize microservice call graphs from Alibaba cluster data, categorizing calling dependencies into three distinct types. Parslo (Mirhosseini et al., 2021) introduces a gradient descent method for decomposing end-to-end SLO budgets into node-specific latency targets. These solutions, however, are computationally intensive and unsuitable for handling dynamically changing request patterns efficiently.
2.3. Network Dynamics Emulation
@@
Networking dynamics significantly influence microservice performance in cloud-edge environments. THUNDERSTORM (Liechti et al., 2019) provides tools for evaluating distributed system performance under dynamic network topologies. Kollaps (Gouveia et al., 2020) introduces a decentralized emulator to simulate dynamic network conditions using simplified topologies. SplayNet (Schmid et al., 2016) extends Splay (Leonini et al., 2009) for arbitrary topology emulation via graph analysis. However, these solutions lack efficient measurement and controllable injection mechanisms essential for accurately evaluating scheduling policies. Existing delay injection methods (Qiu et al., 2020; Wang et al., 2021; Xu et al., 2022b; Fu et al., 2021b, a) typically use static or uniform delay matrices, limiting adaptability and complicating integration into dynamic cloud-edge scenarios.
2.4. Network-aware Microservice Scheduling
@@
Network-aware scheduling strategies such as NetMARKS (Wojciechowski et al., 2021), Marchese et al. (Marchese and Tomarchio, 2022a), and OptTraffic (Zhu et al., 2023) utilize dynamic metrics and infrastructure network conditions to schedule Kubernetes pods optimally. However, these methods insufficiently handle detailed bi-directional traffic metrics and varying cross-node communication latencies, which are critical in real-world cloud-edge deployments.
3. Background and Motivation
3.1. Background
@@
Microservices architecture breaks applications into small, independent services that communicate over networks (Luo et al., 2021; Gan et al., 2019). In cloud-edge computing, these services can run in central cloud data centers, on edge nodes (near end-users/devices), or across both. In practice, an application might be split so that some microservices run in the cloud (for centralized tasks or global data) and others run at the edge (for local data processing or quick response), communicating over the network. This hybrid deployment improves application performance by reducing the distance data must travel, lowering user-perceived latency and offloading work from the cloud to local nodes. However, despite the promising benefits of deploying microservice applications in cloud-edge clusters, there are notable challenges when the deployment goes to a practical scenario.
@@
Figure 1. An envisioned workflow illustrating containerized microservice execution and communication under networking dynamics in a cloud-edge continuum.
3.2. Motivation
3.2.1. Dynamic Networking Conditions
@@
In cloud-edge scenarios, microservice applications frequently experience dynamic and heterogeneous networking conditions. Figure 1 illustrates the envisioned workflow of microservices executing across multiple nodes with fluctuating cross-node latencies and bandwidth availability. Microservices heavily depend on network communications through Remote Procedure Calls (RPCs) or REST APIs, resulting in greater sensitivity to cross-node communication variations. Consequently, fluctuating cross-node latencies and bandwidth constraints significantly affect both end-to-end response times and throughput, making it critical to accurately evaluate how scheduling policies perform under such conditions.
@@
Figure 2. Different request types triggering distinct call-graph structures (application source (Gan et al., 2019)).
3.2.2. Variability in Call-graph Structures
@@
Microservice requests traverse complex call-graphs, which vary significantly depending on the type of incoming requests. As depicted in Figure 2, widely-used benchmark applications such as Social Network (Gan et al., 2019) feature different request types (e.g., compose-post, read-home-timeline, read-user-timeline), each triggering unique call-graph structures. Understanding these variations is crucial, as scheduling policies must handle not only diverse call-graphs but also dynamically changing proportions of request loads. Consequently, evaluating scheduling strategies requires capturing these variations in call-graph structures and their respective workloads accurately.
3.2.3. Imbalanced Traffic Distribution
@@
Figure 3. Imbalanced traffic loads among UM-DM pairs under high workload (5k Queries Per Second) scenarios.
@@
Figure 4. The proposed framework iDynamics for managing different dynamics in cloud-edge continuum.
4. Framework of iDynamics
Motivated by the challenges discussed previously, we propose iDynamics, a unified evaluation framework for implementing and evaluating different scheduling policies of containerized microservices in cloud-edge computing environments. To enable robust policy evaluation, iDynamics comprehensively manages various dynamic conditions, including networking variations, changing call-graph structures, and imbalanced traffic distributions. Figure 4 illustrates the framework’s main components and their interactions during the evaluation and scheduling processes. Key components of iDynamics include:
  • Graph Dynamics Analyzer: Consisting of the UM-DM Traffic Profiler and Call-graph Builder, this component analyzes dynamic call-graphs triggered by different request types and quantifies traffic imbalances under varying workloads (Section § 5).
  • Networking Dynamics Manager: Including an Emulator and a distributed Measurer to manage and emulate realistic networking conditions such as cross-node latency and bandwidth. The Emulator generates configurable dynamics, while the distributed Measurer collects actual communication metrics via daemon agents deployed across cloud-edge nodes (Section § 6).
  • Scheduling Policy Extender: Including a Policy Customization Interface and a Utility Function Module, this component enables rapid development, customization, and evaluation of scheduling strategies. Users can implement tailored scheduling logic and evaluate their policies across diverse dynamic scenarios (Section § 7).
@@ 4. Framework of iDynamics
    SLA violations for policy evaluation are induced by either tightening SLA criteria or injecting intensified dynamics (e.g., increased latency or workload), thus simulating performance degradation. Scheduling decisions are temporarily stored in a decision queue, then filtered based on compliance requirements or migration constraints, enhancing flexibility and control.
@@
5. Design of Graph Dynamics Analyzer
@@
Considering the characteristics of microservice call-graphs described in Section § 3.2 Motivation, the Graph Dynamics Analyzer is designed primarily to analyze topology dynamics (variations in triggered call-graphs) and traffic dynamics (imbalanced traffic distribution) of microservice applications deployed in the cloud-edge continuum. To effectively model these dynamic call-graph structures and traffic distributions, we leverage a service mesh to implement the UM-DM Traffic Profiler and Call-graph Builder as core modules within the Graph Dynamics Analyzer, as shown in Figure  4.
5.1. Service Mesh
@@
A service mesh facilitates real-time analysis and bidirectional monitoring of traffic between multiple upstream microservice (UM) and downstream microservice (DM) replicas. A critical requirement in developing the Graph Dynamics Analyzer component for iDynamics, as shown in Figure 4, is efficiently capturing and analyzing bidirectional traffic between dependent microservices. Collecting metrics is straightforward when each microservice has a single instance; metrics can simply be gathered from the Linux proc file system. However, multiple replicas of upstream or downstream microservices significantly complicate traffic analysis. Therefore, a service mesh is introduced to efficiently observe and manage these complex, multi-instance traffic flows.
5.1.1. Istio Service Mesh Implementation

### v3

# iDynamics: A Novel Framework for Evaluating Microservice Scheduling Policies under Controllable Dynamics in Cloud-Edge Continuum

@@
Abstract
Designing and evaluating microservice scheduling policies is challenging, particularly under dynamic conditions such as complex call-graph dependencies and varying cross-node networking conditions. Moreover, deploying such systems in real-world edge–cloud environments to evaluate scheduling strategies is often impractical due to complexity, cost, and limited accessibility. This highlights the need for an emulation framework that can faithfully emulate the characteristics of the edge–cloud continuum. These characteristics include dynamic topology changes, latency-sensitive service chains, and varying networking conditions, all of which must be accurately modeled for meaningful evaluation.
@@ Abstract
In this work, iDynamics addresses these challenges by providing a configurable and extensible framework that captures the essential dynamics of running microservice applications in cloud-edge environments, enabling systematic development and testing of microservice scheduling strategies. The framework comprises modular components, such as the Graph Dynamics Analyzer, Networking Dynamics Manager, and Scheduling Policy Extender. This enables fine-grained environmental control and facilitates systematic comparisons of different scheduling strategies. Extensive experiments on a real cloud-edge testbed demonstrate that iDynamics effectively captures diverse dynamic scenarios encountered in microservice deployments, offering a robust solution for designing and evaluating different policies under realistic and controllable conditions.
Index Terms:
@@
Microservices, Policy Evaluation, Controllable Dynamics, Cloud-Edge Continuum.
publicationid: pubid: $0$–$00$ 2021 IEEE
I Introduction
@@
Microservices have emerged as an essential architecture in the cloud-edge computing continuum, revolutionizing the design and implementation of distributed applications. Characterized by their modularity and decentralization, microservices enhance flexibility and scalability, effectively meeting dynamic application demands [1, 2, 3]. However, this architectural transition introduces significant challenges, especially for managing complex workflows and intricate service dependencies [4, 5, 6, 7]. Unpredictable user requests, varying execution paths across services, and fluctuating network conditions exacerbate the difficulty of ensuring SLA (Service Level Agreement) compliance for applications deployed across cluster nodes.
@@
Within the cloud-edge continuum, numerous microservices interact frequently and are often shared among multiple applications [8, 9]. Such interactions can lead to performance degradation due to resource contention and cascading delays along request execution paths, directly impacting critical SLA-related metrics such as latency and throughput.
As these challenges arise from various dynamics at different levels, including workload, application, and infrastructure levels, fine-grained scheduling policies and the corresponding evaluation framework should be proposed and evaluated. With the availability of such a policy evaluation framework, it would be more accurate and easier for the cloud practitioners and researchers to accommodate different applications (i.e., microservices) to meet different SLAs from users.
@@
Although various frameworks have been proposed for evaluating scheduling policies for cloud microservices, existing solutions still have limitations. For instance, AutoARMOR [10] automates inter-service access control policy generation but overlooks dynamic networking conditions and is limited in supporting diverse SLA evaluations. OptTraffic [11] optimizes containerized microservice traffic but neglects cross-node latency fluctuations and introduces complexity due to pairwise analysis of dependent container replicas. Similarly, NetMARKS [12] utilizes dynamic network metrics from Istio Service Mesh for pod scheduling, yet it lacks comprehensive bi-directional traffic analysis and can lead to imbalanced load distributions. Other existing methods [13, 14, 15, 16, 17] also face limitations regarding handling call-graph dynamics and networking variations.
@@
To address these gaps, we propose iDynamics, a novel framework that systematically manages multiple layers of dynamics, enabling precise evaluation of scheduling policies under realistic conditions. With iDynamics, cloud practitioners and researchers can observe the impact of dynamics on application performance and identify SLA violations effectively. Figure 4 illustrates the high-level design of our proposed framework, highlighting its modular components. To our knowledge, iDynamics is the first unified evaluation framework specifically designed with controllable dynamics for microservice applications in cloud-edge continuum environments. Our primary contributions are:
@@
The rest of the paper is structured as follows: Section 2 reviews related work. Section 3 presents the background and challenges. Section 4 introduces the iDynamics framework and main components. Section 5 describes the design of the Graph Dynamics Analyzer, and Section 6 details the Networking Dynamics Manager. Section 7 explains the Scheduling Policy Extender. Section 8 provides performance evaluations and detailed case studies. Finally, Section 9 concludes the paper and outlines future work.
TABLE I: Comparison of representative work.
Work Microservice Management Call-graph Dynamics Analysis Dynamic Networking Control Scheduling Policy Extender
SLA Guarantees Scheduling E2E Perf. Dependency Traffic Analysis Emulation Delays Bandwidth Measurement Extensible Interfaces Utility Modules Cloud-Edge
FIRM [18]
Erms [9]
GandSLAm [19]
Sage [20]
Kollaps [21]
NetMARKS [12]
OptTraffic [11]
iDynamics (This Work)
II Related Work
@@
The management of microservices in different scenarios has been extensively explored. This section reviews existing literature emphasizing microservice management, call-graph analysis, network dynamics emulation, and network-aware scheduling. Besides, Table I compares iDynamics with representative work from different aspects.
II-A Microservice Management
@@
FIRM [18] leverages telemetry metrics and machine learning models to fine-tune microservice management by identifying SLO violations and resource contentions, subsequently mitigating violations through reprovisioning actions. Erms [9] formulates resource scaling models specifically addressing latency goals for shared microservices with extensive call graphs. GrandSLAm [19] predicts microservice execution stages’ completion times, dynamically batching and reordering requests based on these predictions. Nevertheless, these approaches often neglect dynamic cross-node delays and the impacts of bi-directional communication between replicated microservices.
II-B Call-graph Dynamics Analysis
@@
Sage [20] employs graphical Bayesian models for root cause analysis of cascading QoS violations in interactive microservices, prioritizing practicality and scalability. Tian et al. [22] develop a synthetic workload generator for DAG jobs by analyzing large-scale cluster traces. Luo et al. [23] characterize microservice call graphs from Alibaba cluster data, categorizing calling dependencies into three distinct types. Parslo [24] introduces a gradient descent method for decomposing end-to-end SLO budgets into node-specific latency targets. These solutions, however, are computationally intensive and unsuitable for handling dynamically changing request patterns efficiently.
II-C Network Dynamics Emulation
@@
Networking dynamics significantly influence microservice performance in cloud-edge environments. THUNDERSTORM [25] provides tools for evaluating distributed system performance under dynamic network topologies. Kollaps [21] introduces a decentralized emulator to simulate dynamic network conditions using simplified topologies. SplayNet [26] extends Splay [27] for arbitrary topology emulation via graph analysis. However, these solutions lack efficient measurement and controllable injection mechanisms for accurately evaluating scheduling policies. Existing delay injection methods [18, 28, 29, 30, 31] typically use static or uniform delay matrices, limiting adaptability and complicating integration into dynamic cloud-edge scenarios.
II-D Network-aware Microservice Scheduling
@@
Network-aware scheduling strategies such as NetMARKS [12], Marchese et al. [17], and OptTraffic [11] utilize dynamic metrics and infrastructure network conditions to schedule Kubernetes pods optimally. However, these methods are insufficiently handling detailed bi-directional traffic metrics and varying cross-node communication latencies, which are critical in real-world cloud-edge deployments.
III Background and Challenges
III-A Background
@@
Microservices architecture breaks applications into small, independent services that communicate over networks [23, 32]. In cloud-edge computing, these services can run in central cloud data centers, on edge nodes (near end-users/devices), or across both. In practice, an application might be split so that some microservices run in the cloud (for centralized tasks or global data) and others run at the edge (for local data processing or quick response), communicating over the network. This hybrid deployment improves application performance by reducing the distance data must travel, lowering user-perceived latency, and offloading work from the cloud to local nodes. However, despite the promising benefits of deploying microservice applications in cloud-edge clusters, there are notable challenges when the deployment goes to a practical scenario.
@@
Figure 1: An envisioned workflow illustrating containerized microservice execution and communication under networking dynamics in a cloud-edge continuum.
III-B Challenges
III-B1 Dynamic Networking Conditions
@@
In cloud-edge scenarios, microservice applications frequently experience dynamic and heterogeneous networking conditions. Figure 1 illustrates the envisioned workflow of microservices executing across multiple nodes with fluctuating cross-node latencies and bandwidth availability. Microservices heavily depend on network communications through Remote Procedure Calls (RPCs) or REST APIs, resulting in greater sensitivity to cross-node communication variations. Consequently, fluctuating cross-node latencies and bandwidth constraints significantly affect end-to-end response times and throughput, making it critical to evaluate how scheduling policies perform under such conditions accurately.
@@
Figure 2: Different request types triggering distinct call-graph structures (application source [32]).
Figure 3: Imbalanced traffic loads among UM-DM pairs under high workload (5k Queries Per Second) scenarios.
III-B2 Variability in Call-graph Structures
@@
Microservice requests traverse complex call graphs, which vary significantly depending on the type of incoming requests. As depicted in Figure 2, widely-used benchmark applications such as Social Network [32] feature different request types (e.g., compose-post, read-home-timeline, read-user-timeline), each triggering unique call-graph structures. Understanding these variations is crucial, as scheduling policies must handle diverse call-graphs and dynamically changing proportions of request loads. Consequently, evaluating scheduling strategies requires accurately capturing these variations in call-graph structures and their respective workloads.
III-B3 Imbalanced Traffic Distribution
@@
Figure 4: The proposed iDynamics for managing different dynamics and implementing scheduling policies to mitigate SLA violations in cloud-edge continuum.
IV Framework of iDynamics
Motivated by the challenges discussed previously, we propose iDynamics, a unified evaluation framework for implementing and evaluating different scheduling policies of containerized microservices in cloud-edge computing environments. To enable robust policy evaluation, iDynamics comprehensively manages various dynamic conditions, including networking variations, changing call-graph structures, and imbalanced traffic distributions. Figure 4 illustrates the framework’s main components and interactions during the evaluation and scheduling processes. Key components of iDynamics include:
  • Graph Dynamics Analyzer: Consisting of the UM-DM Traffic Profiler and Call-graph Builder, this component analyzes dynamic call-graphs triggered by different request types and quantifies traffic imbalances under varying workloads (Section § V).
  • Networking Dynamics Manager: Including an Emulator and a distributed Measurer to manage and emulate realistic networking conditions such as cross-node latency and bandwidth. The Emulator generates configurable dynamics, while the distributed Measurer collects actual communication metrics via daemon agents deployed across cloud-edge nodes (Section § VI).
  • Scheduling Policy Extender: Including a Policy Customization Interface and a Utility Function Module, this component enables rapid development, customization, and evaluation of scheduling strategies. Cloud practitioners and researchers can implement tailored scheduling logic and evaluate their policies across diverse dynamic scenarios (Section § VII).
@@ IV Framework of iDynamics
    SLA violations for policy evaluation are induced by either tightening SLA criteria or injecting intensified dynamics (e.g., increased latency or workload), thus causing performance degradation. Scheduling decisions are temporarily stored in a decision queue, then filtered based on compliance requirements or service migration constraints, enhancing flexibility and control.
@@
V Design of Graph Dynamics Analyzer
@@
Considering the characteristics of microservice call-graphs described in Section § III-B Challenges, the Graph Dynamics Analyzer is designed primarily to analyze topology dynamics (variations in triggered call-graphs) and traffic dynamics (imbalanced traffic distribution) of microservice applications deployed in the cloud-edge continuum. To effectively model these dynamic call-graph structures and traffic distributions, we leverage a service mesh to implement the UM-DM Traffic Profiler and Call-graph Builder as core modules within the Graph Dynamics Analyzer, as shown in Figure  4.
V-A Service Mesh
@@
A service mesh facilitates real-time analysis and bidirectional traffic monitoring between multiple upstream microservices (UM) and downstream microservices (DM) replicas. A critical requirement in developing the Graph Dynamics Analyzer component for iDynamics, as shown in Figure 4, is efficiently capturing and analyzing bidirectional traffic between dependent microservices. Collecting metrics is straightforward when each microservice has a single instance; metrics can be gathered from the Linux proc file system. However, multiple replicas of upstream or downstream microservices significantly complicate traffic analysis. Therefore, a service mesh is introduced to observe and manage these complex, multi-instance traffic flows efficiently.
V-A1 Istio Service Mesh Implementation